Circular No. 8058

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003B, 2003Q, 2003R, 2003S, AND 2003T Several additional supernova discoveries on unfiltered CCD images

have been reported, with the tabulated magnitudes below all stated to be approximate. SNe 2003Q and 2003R were reported by D. Poznanski, A. Gal-Yam, K. Sharon, and D. Maoz, Tel Aviv University; and B. T. Jannuzi, National Optical Astronomy Observatory (NOAO), from images of the NOAO Deep Wide Field Survey field J0205–0455 obtained by Jannuzi with the 4-m Mayall telescope (+ Mosaic Camera) at Kitt Peak. SNe 2003S and 2003T were found via LOTOSS (cf. IAUC 7906) — the former reported by M. Papenkova and W. Li (KAIT discovery) and the latter by M. Schwartz and Li (Tenagra II discovery).

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003Q Jan. 2  2^{h}05^{m}47.^{s}55 − 4^{o}52′48′′.0 20  —
2003R Jan. 2  2 05 11.27 − 4 42 30.0 22  —
2003S Jan. 24.6 13 41 19.49 +55 40 36.5 18.0 1′′.1 W, 3′′.3 S
2003T Jan. 25.4 9 14 11.06 +16 44 48.0 18.1 25′′.0 W, 14′′.9 N

Additional magnitudes: SN 2003Q, 2002 Oct. 11, [26: (Jannuzi). SN 2003R, 2002 Oct. 11, [26: (Jannuzi). SN 2003S in MCG +09-22-94, 2002 June 14.3 UT, [19.0: (KAIT); 2003 Jan. 26.5, 18.0: (Tenagra II). SN 2003T in UGC 4864, Jan. 5.4, [19.0: (KAIT); 27.4, 17.9: (Tenagra II).

A. V. Filippenko and R. Chornock, University of California, Berkeley, report that inspection of CCD spectra (range 310–1000 nm), obtained on Jan. 7 UT with the Keck I 10-m telescope (+ LRIS), shows that SN 2003Q is of type Ia, probably within a week of maximum brightness; the approximate redshift is 0.21, based on a comparison with the spectrum of SN 1989B. SN 2003R is also of type Ia, with a redshift of 0.668 (measured from narrow [O II] emission), but the S/N ratio is fairly low.

T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2003S, obtained by P. Berlind on Jan. 27.48 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova near maximum. Narrow emission lines superposed on the spectrum of the supernova indicate a recession velocity of 11700 km/s for the host galaxy. Using this value, the supernova expansion velocity is ∼ 10600 km/s for Si II (rest 635.5 nm).

Prediscovery unfiltered CCD images taken on 2002 Dec. 28.09 UT by M. Motta at Center Harbor, NH, show SN 2003B at mag 16.4.

2003 January 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green